Given a number such that the product of 2 more than the number times 3 less than the same number is 44. Find the number
Answer by ankor@dixie-net.com(22740)

Given a number such that the product of 2 more than the number times 3 less than the same number is 44.
Find the number
:
(n+2)*(n-3) = 44
FOIL
n^2 - 3n + 2n - 6 = 44
combine on the left to form a quadratic equation
n^2 - n - 6 - 44 = 0
n^2 - n - 50 = 0
unfortunately, this will not factor, (solution is not an integer)
Use the quadratic formula and the approx solutions are
n = +7.59
n = -6.59